What is the answer to: Guess my number: One number is five more than a second number. The product of the two numbers is 3300. Find the two numbers.
Answer by algebrahouse.com(1659)

"What is the answer to: Guess my number: One number is five more than a second number. The product of the two numbers is 3300. Find the two numbers."

x = one number
x + 5 = other number {one number is five more than the other}

x(x + 5) = 3300 {their product is 3300}
x² + 5x = 3300 {used distributive property}
x² + 5x - 3300 = 0 {subtracted 3300 from both sides}
(x + 60)(x - 55) = 0 {factored into two binomials}
x + 60 = 0 or x - 55 = 0 {set each factor equal to 0}
x = -60 or x = 55 {subtracted 60 and added 55, respectively}
x + 5 = -55 or x + 5 = 60 {substituted -60 and 55, in for x, into x + 5}

-60 and -55 are the two numbers
or
55 and 60 are the two numbers
www.algebrahouse.com